Please I need help with this one. Thanks million.
Solve by the elimination method.
7r+3s=38
3r+7s=66
--------------
21r+9s= -114
-21r-49= -462
-58= -576
S=? 9.9....?
--------------
7r-3(3)=-38
R=?

Answer by scott8148(6628)

7r+3s=38
3r+7s=66
--------------

21r+9s= -114 ___ why is the 114 negative?
try ___ 21r + 9s = 114

-21r-49= -462 ___ what happened to the "s"?
try ___ -21r - 49s = -462

-58= -576 ___ -40s = -348 ___ s = 8.7
S=? 9.9....?
--------------

7r + 3 (8.7) = 38 ___ 7r + 26.1 = 38 ___ 7r = 11.9 ___ r = 1.7
R=?